What is equilibrium? Suppose, demand function for Singara is QD = 50 – 2P and supply
function is QS= 20 + 3P. Find out -
(i) The equilibrium price and quantity of Singara.
(ii) If the supply increases to QS = 30+3P, what will happen to the equilibrium
price and quantity?
(iii) If government imposes a price ceiling at Tk. 3, what will happen to the market?
1
Expert's answer
2019-10-03T09:19:12-0400

i)At equilibrium price quantity demanded =quantity supplied


"50-2p=20+3p"


"5p=30""p=Tk.6"

Equilibrium quantity=

"50-2\times6=38"

ii)

"50-2p=30+3p"


"20=5p,p=Tk.4"

Equilibrium quantity =

"50-2\times4=42"

Equilibrium price falls to Tk. 4 while equilibrium quantity increases to 42

iii)Quantity demanded is

"50-2\times3=44"

Quantity supplied is

"20+3\times3=29"

The price ceiling will create a shortage of 15 units in the market